### v1

@@ IV-D Performance Analysis over Varying Reverberation Times and SNR Levels
Here, we evaluate the performance of both the proposed method and the baseline method for varying reverberation times $T_{60}$ and varying SNRs, averaging within the spherical sweet area at the bandwidth for $15$ time frames. The evaluation is based on corresponding sound pressures at $107$ observation points as Sec. IV-C. The estimated ReHCs are used in the proposed method as Sec. IV-A.
@@
Firstly, we evaluate the performance of these two methods over varying reverberation times $T_{60}$ as shown in TABLE I. Here, we vary the $T_{60}$ from 00 s to $0.4$ s, while the SNR is set to be 00 dB. With the reverberation time $T_{60}$ increased, both methods achieved higher estimation error, lower SDR, and lower NR. In addition, the proposed method always achieved lower estimation error, higher SDR, and comparable NR than the baseline method for varying $T_{60}$ (more than $25$ dB NR is effective enough).
@@
Secondly, we evaluate the performance of these two methods over varying SNR levels as shown in TABLE II. Here, we vary the SNR level from $5$ dB to $-5$ dB, and remain the $T_{60}$ as $0.2$ s. TABLE II shows that the proposed method maintained stable performance for varying SNR levels in terms of estimation error (about $-19$ dB) and SDR (about $22$ dB), always outperforming the baseline method. By contrast, with the SNR level decreased, both estimation error and SDR achieved by the baseline method declined. In addition, both the proposed method and the baseline method achieved comparable and stable NR (more than about $28$ dB) for varying SNR levels.
V Conclusion
@@
In this paper, we propose a spherical harmonic (SH)-domain minimum variance distortionless response (MVDR) method to estimate the desired sound field from the mixture recording at a spherical microphone array. We use a cost function with a set of spatial constraints to extract desired SH coefficients and suppress SH coefficients of interference and sensor noises. The field due to the desired sound source within the sweet area can be hence reconstructed. Simulation results show that the proposed method can outperform the baseline methods within the sweet area in a reverberant room. In the future, we plan to further examine the proposed method in real-world scenarios.

### v2

@@ 4.4 Performance Analysis over Varying Reverberation Times and SNR Levels
Here, we evaluate the performance of both the proposed method with estimated ReHCs and the baseline method for varying reverberation times $T_{60}$ and varying SNRs, averaging within the spherical sweet area at the bandwidth for $15$ time frames. As Sec. 4.3, three indexes are evaluated at these $107$ observation points.
@@
Firstly, we evaluate the performance of these two methods over varying reverberation times $T_{60}$ as shown in TABLE 1. Here, we vary the $T_{60}$ from 0 s to $0.4$ s, while the SNR is set to be 0 dB. With the reverberation time $T_{60}$ increased, both methods achieved higher estimation error, lower SDR, and lower NR. In addition, the proposed method always achieved lower estimation error, higher SDR, and comparable NR than the baseline method for varying $T_{60}$ (more than $25$ dB NR is effective enough).
@@
Secondly, we evaluate the performance of these two methods over varying SNR levels as shown in TABLE 2. Here, we vary the SNR level from $5$ dB to 5-5 dB, and remain the $T_{60}$ as $0.2$ s. TABLE 2 shows that the proposed method maintained stable performance for varying SNR levels in terms of estimation error (about 19-19 dB) and SDR (about $22$ dB), always outperforming the baseline method. By contrast, with the SNR level decreased, both estimation error and SDR achieved by the baseline method declined. In addition, both the proposed method and the baseline method achieved comparable and stable NR (more than about $28$ dB) for varying SNR levels.
5 Conclusion
@@
In this paper, we propose a spherical harmonic (SH)-domain minimum variance distortionless response (MVDR) method to estimate the desired sound field from the mixture recording at a spherical microphone array. We use a cost function with a set of spatial constraints to extract desired SH coefficients and suppress SH coefficients of interference and sensor noises. The field due to the desired sound source within the sweet area can be hence reconstructed. Simulation results show that the proposed method can outperform the baseline methods within the sweet area in a reverberant room. In the future, we plan to theoretically analyze the influence of ReHC accuracy on the performance of our proposed method.